OpenStudy (anonymous):

given 2x-y=6, solve for y

OpenStudy (johnweldon1993):

so you would want to isolate 'y' on the left side of the equation...how would we do that?

OpenStudy (anonymous):

-2x-6

OpenStudy (anonymous):

Thankyou !

OpenStudy (anonymous):

just kidding its 2x-6

OpenStudy (johnweldon1993):

@connor.sheedy please refrain from giving the answer...especially if it is incorrect 2x - y = 6 -y = 6 - 2x y = -6 + 2x y = 2x - 6